OpenStudy (call_me_abby):

How to solve 192 = x^2 + 12x

OpenStudy (kittyloli):

You have to get x by its self... \[192=x ^{2}+12x\] then \[x ^{2}\] Counts as 1 x, so \[x ^{2}+12x= 13x ^{2}\] \[192=13x ^{2}\] Now, to get rid of the \[x ^{2}\] You need to square root it, so \[\sqrt{192}=\sqrt{13x}=13.856\] Round to the nearest tenth 13.9=13x Then divide \[13.9\div13x=1.069\] Round to the nearest tenth x=1.1

OpenStudy (kittyloli):

Hopefully I did that correctly

OpenStudy (wolf1728):

No that's not the answer I got

OpenStudy (dustin_whitelock):

That's wrong @Kittyloli You cannot add x^2 and x, both are different.

OpenStudy (kittyloli):

HECK

OpenStudy (kittyloli):

This is why I can't do math

OpenStudy (zzr0ck3r):

\(x^2+12x-192=0\) Can you think of two numbers that multiply to -192 but add to 12?

OpenStudy (wolf1728):

192 = x^2 + 12x Rewrite the equation as x^2 +12x -192 = In this equation a=1 b=12 and c=-192 Do you know the quadratic formula?

OpenStudy (call_me_abby):

oh I do! Thank you so much I think I got it now

OpenStudy (wolf1728):

Wow okay Abby - that was easy!!

OpenStudy (dustin_whitelock):

Try factorization This rule is called middle term factorization factorize 192 You get 192 = 2*2*2*2*2*2*8 Now arrange 192 as sum or subtraction of the terms to btain 12 You see 2*2*2*3-2*2*2=24-8=12 Now 12x= 24x-8x Now, \[x ^{2} + 12x-192=0\] or, \[x ^{2}+ 24x-8x-192=0\] or, x(x+4)-8(x+24)=0 or, (x+24) (x-8)=0 or, x=-24,8 @wolf1728, I don't think quadratic equation is an intelligent way to answer this simple question. Use that rule when you cannot factorize easily.

OpenStudy (call_me_abby):

@Dustin_Whitelock 24x — 8x equals 16x though

OpenStudy (dustin_whitelock):

@Call_Me_Abby Congratulations, for finding out the intentional mistake. Now since you know how to obtain the answer, do the factorization correctly and get your answer.

OpenStudy (call_me_abby):

Yep I did both your way and the quadratic formula way and got my answer

OpenStudy (dustin_whitelock):

:)

OpenStudy (wolf1728):

24 and 8 are NOT the answers 192 factors into 2*2*2*2*3*19 x^2 +12x -192 a=1 b=12 and c=-192 Using quadratic formula: x = -12 +- sqroot (144 -4*1*-192) / 2*1 x = -12 +- sqroot (144 + 768) / 2 x = -12 +-(30.1993377411)/2 x1= (-12 + 30.1993377411)/2 x1 = 9.0996688706 x2= (-12 - 30.1993377411)/2 x2 = -21.0996688706

OpenStudy (call_me_abby):

I got 9.1, Dustin_Whitelock made the mistake on purpose so they wouldn't be giving out the answer

OpenStudy (wolf1728):

Okay Abby well I gave out the answers to many decimal places. (Mainly because I thought the equation was already solved.) Anyway 9.1 is one answer and you can see the other one.

OpenStudy (call_me_abby):

Thankss now I know for sure for sure that the answer's right

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

The actual best method here is completing the square. The middle term is even and the leading coefficient is 1. This is fact. jk

OpenStudy (mathstudent55):

\(192 = x^2 + 12x\) \(x^2 + 12x - 192 = 0\) Use the quadratic formula: \(\large x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2b} \) \(\large x = \dfrac{-12 \pm \sqrt{(-12)^2 - 4(1)(-192)}}{2(1)} \) \(\large x = \dfrac{-12 \pm \sqrt{144 + 768}}{2} \) \(\large x = \dfrac{-12 \pm \sqrt{912}}{2} \) \(\large x = \dfrac{-12 \pm 4\sqrt{57}}{2} \) \(\large x = -6 \pm 2\sqrt{57} \)